# Improving Genomic Prediction using High-dimensional Secondary Phenotypes: the Genetic Latent Factor Approach

### v1

@@ 2.7. Estimation of factor scores (step VI)
One of the characteristics of Thomson regression that we maintain in our modified approach is the orthogonality of the factor scores. We have thus reduced the possibly high-dimensional and collinear secondary data to a much smaller set of orthogonal factors, without losing interpretability. These orthogonal factors can easily and effectively be used in any multivariate genomic prediction approach. We opt to include (a subset of) the $\tilde{m}$ factors, the relevant covariance matrices, and the marker-based kinship matrix in the standard multivariate BLUP equations as described in the following section, thus finally integrating genomic and secondary data.
2.8. Factor selection and multivariate genomic prediction (step VII)
Having reduced the dimensionality of the secondary data, we replace the original features in the plot-level data matrix (1) by their projections:
$\mathbf{Y}_{\tilde{\Xi}f}=\begin{bmatrix}\tilde{\bm{\Xi}}&\mathbf{y}_{f}\end{bmatrix}\text{,}$
@@ 2.8. Factor selection and multivariate genomic prediction (step VII)
and use sums of squares to obtain estimates for the genetic and residual covariance matrices of the factors and focal trait. We then convert the plot-level factor scores to BLUEs $\bar{\mathbf{Y}}_{\tilde{\Xi}f}$ by taking the genotypic means. Note that we have not used the focal trait data $\mathbf{y}_{f}$ in any of the previous six steps detailed above. As a result, not all of the $\tilde{m}$ factors may be relevant to the multivariate genomic prediction problem. As a form of subset selection we perform an exhaustive search for the best subset of $\tilde{m}^{\star}\leq\tilde{m}$ factors for predicting the focal trait BLUEs using the factor BLUEs in linear regression, using adjusted $R^{2}$ as the accuracy measure [32]. Following this search, we subset the BLUEs to obtain $\bar{\mathbf{Y}}_{\tilde{\Xi}^{\star}f}\in\mathbb{R}^{n\times(\tilde{m}^{\star}+1)}$ subset the covariance matrices of factors and focal trait accordingly:
@@
We finally divide $\hat{\bm{\Sigma}}^{\epsilon\star}$ by the number of replicates and combine secondary and genomic data using the standard BLUP equations (see SM Section S.2).
The integration of secondary and genomic data thus follows from the use of the marker-based kinship matrix and the inclusion of the factor scores in the data matrices, along with the addition of the factors to the genetic and residual covariance matrices. If secondary data is only available for the training set (CV1), the extra features refine estimates of genetic (i.e., marker) effects for training genotypes. These improved training set estimates are then translated to test set estimates. In CV2, where secondary data is available for the test set too, the extra features provide direct information on the genetic effects for the test set, improving estimates of the focal trait directly. To evaluate the BLUP equations described in SM Section S.2 we use a fast algorithm that decreases the complexity of these calculations from $O(n_{g}^{3}(\tilde{m}+1)^{3})$ to $O(n_{g}(\tilde{m}+1)^{2}+(\tilde{m}+1)^{3})$ [33].
On a final note, consider that as the dimensionality reduction described in steps I to VI is entirely unsupervised, gfBLUP is flexible with regards to what secondary feature data is used. If secondary features are only available for the training set (CV1), estimation of the covariance matrices, redundancy filtering, regularization, estimation of $\bm{\Theta}$, etc. is based only on training data and factor scores are only obtained for the training set. In CV2, where secondary features are also available for the test set, this data can be included effortlessly as well using the multivariate BLUP equations found in SM Section S.2.
@@ 3. Simulations
We compared gfBLUP to a number of alternative approaches (siBLUP, lsBLUP, MegaLMM, deep learning (multiMLP: multivariate multilayer perceptron), univariate gBLUP, see SM Section S.5 for details on these methods) in terms of performance in different high-dimensional data settings. The following two sections provide details on the simulation process as well as the results.
@@ 3.1. Setup
Figure 2. Prediction accuracies for several methods in the CV1 and CV2 scenarios as we vary the focal trait communality, focal trait heritability, and secondary feature heritabilities. Details regarding these simulations can be found in Section 3.1.

### v3

@@ 2.2.7. Estimation of factor scores (step VI)
One of the characteristics of Thomson regression that we maintain in our modified approach is the orthogonality of the factor scores. We have thus reduced the possibly high-dimensional and collinear secondary data to a much smaller set of orthogonal factors, without losing interpretability. These orthogonal factors can easily and effectively be used in any multivariate genomic prediction approach. We use the next section to describe the multivariate genomic prediction step using the factor scores, target trait, and marker-based kinship matrix, thus finally integrating genomic and secondary feature data.
2.3. Modeling of latent factors and focal trait
Having reduced the dimensionality of the secondary data we now proceed with the multivariate genomic prediction. The remainder of this section contains a description of the model and assumptions, estimation of covariance matrices, selection of relevant factors, and calculations resulting in the final focal trait predictions.
2.3.1. Model and assumptions
Consider the data matrix $\mathbf{Y}_{\tilde{\Xi}f}\in\mathbb{R}^{n\times(\tilde{m}+1)}$ containing the latent factor scores and focal trait. The matrix $\mathbf{Y}_{\tilde{\Xi}f}$ can be partitioned:
(11) $\mathbf{Y}_{\tilde{\Xi}f}=\begin{bmatrix}\mathbf{Y}_{\tilde{\Xi}}&\mathbf{y}_{f}\end{bmatrix}\text{,}$
where the matrix $\mathbf{Y}_{\tilde{\Xi}}$ is the $n\times\tilde{m}$ matrix containing the plot-level factor scores and $\mathbf{y}_{f}$ is the $n\times 1$ column vector containing the focal trait. Like for the secondary features, these factor scores and focal trait are composed of genetic and residual components:
$\mathbf{Y}_{\tilde{\Xi}f}=\mathbf{G}_{\tilde{\Xi}f}+\mathbf{E}_{\tilde{\Xi}f}\text{.}$
We again assume that the genetic components are independent of the residual components so that
$\operatorname{vec}({\mathbf{Y}_{\tilde{\Xi}f}})=\operatorname{vec}({\mathbf{G}_{\tilde{\Xi}f}})+\operatorname{vec}({\mathbf{E}_{\tilde{\Xi}f}})\sim\mathcal{N}_{np}(\mathbf{0},\bm{\Sigma}^{g}\otimes\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}+\bm{\Sigma}^{\epsilon}\otimes\mathbf{I}_{n})\text{,}$
where $\mathbf{Z}\mathbf{K}\mathbf{Z}^{\top}$ and $\mathbf{I}_{n}$ are the row covariances of $\mathbf{G}_{\tilde{\Xi}f}$ and $\mathbf{E}_{\tilde{\Xi}f}$, respectively. The $n\times n_{g}$ incidence matrix $\mathbf{Z}$ links individuals to genotypes. Similarly, $\bm{\Sigma}^{g}$ and $\bm{\Sigma}^{\epsilon}$ are the column covariances of $\mathbf{G}_{\tilde{\Xi}f}$ and $\mathbf{E}_{\tilde{\Xi}f}$. These covariance matrices can be block-partitioned:
$\begin{split}\bm{\Sigma}^{g}&=\begin{pmatrix}\bm{\Sigma}_{\tilde{\Xi}\tilde{\Xi}}^{g}&\bm{\sigma}_{\tilde{\Xi}f}^{g}\\ \bm{\sigma}_{f\tilde{\Xi}}^{g}&\sigma_{ff}^{g}\\ \end{pmatrix}=\begin{pmatrix}\bm{\Sigma}_{\tilde{\Xi}.}^{g}\\ \bm{\sigma}_{f.}^{g}\end{pmatrix}\\ \bm{\Sigma}^{\epsilon}&=\begin{pmatrix}\bm{\Sigma}_{\tilde{\Xi}\tilde{\Xi}}^{\epsilon}&\bm{\sigma}_{\tilde{\Xi}f}^{\epsilon}\\ \bm{\sigma}_{f\tilde{\Xi}}^{\epsilon}&\sigma_{ff}^{\epsilon}\\ \end{pmatrix}=\begin{pmatrix}\bm{\Sigma}_{\tilde{\Xi}.}^{\epsilon}\\ \bm{\sigma}_{f.}^{\epsilon}\end{pmatrix}\text{,}\end{split}$
where $\bm{\Sigma}_{\tilde{\Xi}\tilde{\Xi}}^{g}$ and $\bm{\Sigma}_{\tilde{\Xi}\tilde{\Xi}}^{\epsilon}$ contain the genetic and residual covariances among the latent factors present in the sample, $\sigma_{ff}^{g}$ and $\sigma_{ff}^{\epsilon}$ are the scalar focal trait genetic and residual variances, and $\bm{\sigma}_{\tilde{\Xi}f}^{g}=(\bm{\sigma}_{f\textbf{s}}^{g})^{\top}$ and $\bm{\sigma}_{\textbf{s}f}^{\epsilon}=(\bm{\sigma}_{f\textbf{s}}^{\epsilon})^{\top}$ contain the genetic and residual covariances between latent factors and the focal trait. We estimate the above covariance matrices using the plot-level data as described in Section 2.2.2.
2.3.2. Factor selection and multivariate genomic prediction
@@
We now convert the plot-level data to BLUEs $\bar{\mathbf{Y}}_{\tilde{\Xi}f}$ by taking the genotypic means. Note that we have not used the focal trait data $\mathbf{y}_{f}$ in the modeling of the secondary features. As a result, not all of the $\tilde{m}$ factors may be relevant to the multivariate genomic prediction problem. As a basic form of subset selection we perform an exhaustive search for the best subset of $\tilde{m}^{\star}\leq\tilde{m}$ factors for predicting the focal trait BLUEs $\bar{\mathbf{y}}_{f}$ using the factor score BLUEs $\bar{\mathbf{Y}}_{\tilde{\Xi}}$ in multiple linear regression, using adjusted $R^{2}$ as the accuracy measure (Lumley, 2020). Following this search, we subset the BLUEs to obtain $\bar{\mathbf{Y}}_{\tilde{\Xi}^{\star}f}\in\mathbb{R}^{n\times(\tilde{m}^{\star}+1)}$ and subset the estimated covariance matrices of factor scores and focal trait accordingly:
@@
We finally divide $\hat{\bm{\Sigma}}^{\epsilon\star}$ by the number of replicates and combine secondary and genomic data using the standard multivariate BLUP equations (see SM Section 1 and, e.g., (Runcie and Cheng, 2019; Arouisse et al., 2021)).
The integration of secondary and genomic data thus follows from the use of the marker-based kinship matrix and the inclusion of the factor scores as additional traits in the multivariate genomic prediction. If secondary data are only available for the training set (CV1), the extra features refine estimates of genetic (i.e., marker) effects for training genotypes. These improved training set estimates are then translated to test set estimates. In CV2, where secondary data are available for the test set too, the extra features provide direct information on the genetic effects for the test set, improving estimates of the focal trait directly. To evaluate the BLUP equations described in SM Section 1 we use a fast algorithm that decreases the complexity of these calculations from $O(n_{g}^{3}(\tilde{m}+1)^{3})$ to $O(n_{g}(\tilde{m}+1)^{2}+(\tilde{m}+1)^{3})$ (Dahl et al., 2013).
On a final note, consider that as the dimensionality reduction described in Section 2.2 is entirely unsupervised, glfBLUP is flexible with regards to what secondary feature data are used. If secondary features are only available for the training set (CV1), estimation of the covariance matrices, redundancy filtering, regularization, estimation of $\bm{\Theta}$, etc. is based only on training data and factor scores are only obtained for the training set. In CV2, where secondary features are also available for the test set, these data can be included effortlessly as well using the multivariate BLUP equations found in SM Section 1.
@@ 3. Simulations
We compared glfBLUP to a number of alternative approaches in terms of performance in different high-dimensional data settings. We compared glfBLUP to siBLUP, lsBLUP, MegaLMM, deep learning (multiMLP: multivariate multilayer perceptron), and the univariate gBLUP model, see SM Section 4 for details on these methods. The following two subsections provide details on the simulation process as well as the results.
@@ 3.1. Setup
We simulated a range of different high-dimensional datasets to evaluate model performance under varying circumstances (Figure 1). We used $n_{g}=500$ randomly selected genotypes (1500 SNP markers) from the RegMap Arabidopsis thaliana panel (Horton et al., 2012) to construct a kinship matrix and simulated $r=2$ replicates for a total number of $n=n_{g}r=1000$ individuals. Consider the model for the observed secondary features and focal trait:
(12) $\mathbf{Y}=\mathbf{Z}\left[\left(\bm{\Xi}\mathbf{\Lambda}^{\top}+\bm{\Gamma}\right)\mathbf{S}^{g}\right]+\mathbf{E}\mathbf{S}^{\epsilon}\text{,}$
where $\mathbf{Y}\in\mathbb{R}^{n\times(p+1)}$ is the data matrix, $\bm{\Xi}$ is the $n_{g}\times m$ matrix of factor scores, $\mathbf{\Lambda}$ is the $(p+1)\times m$ matrix of loadings, $\bm{\Gamma}\in\mathbb{R}^{n_{g}\times(p+1)}$ is the matrix containing factor model errors, and $\mathbf{E}$ is the $n\times(p+1)$ matrix containing correlation scale residuals. The $n\times n_{g}$ design matrix $\mathbf{Z}$ links records to genotypes. The matrices $\mathbf{S}^{g}$ and $\mathbf{S}^{\epsilon}$ are diagonal matrices containing the genetic and residual standard deviations of the secondary features and focal trait. The correlation scale genetic components of the phenotypes are thus defined as $\mathbf{G}=\bm{\Xi}\mathbf{\Lambda}^{\top}+\bm{\Gamma}$.
To simulate the data we first randomly sampled $m=8$ factor scores for each genotype from a multivariate normal distribution taking relationships between genotypes as defined in the kinship into account:
(13) $\operatorname{vec}(\bm{\Xi})\sim\mathcal{N}_{n_{g}m}\left(\mathbf{0},\mathbf{\Phi}\otimes\mathbf{K}\right)\text{,}$
where $\mathbf{\Phi}=\mathbf{I}_{m}$, thus assuming that factors are independent. Next we sampled the factor model errors:
(14) $\operatorname{vec}(\bm{\Gamma})\sim\mathcal{N}_{n_{g}p}\left(\mathbf{0},\mathbf{\Psi}\otimes\mathbf{K}\right)\text{,}$
[…]
We constructed the loadings matrix $\mathbf{\Lambda}\in\mathbb{R}^{(p+1)\times m}$ so that eight blocks of $100$ secondary features ($s_{1},\dots,s_{p}$) load on $m=8$ genetic latent factors with a single feature loading on a single factor. Part of the focal trait ($y$) phenotype is generated by $m_{s}=4$ latent signal factors ($LSF$) through loadings. The focal trait does not load on the remaining $m_{n}=4$ factors (latent noise factors, $LNF$), meaning that the $400$ secondary features generated by the $LNFs$ are not genetically correlated to the focal trait. The loading for each of the secondary features $s_{q}$ was obtained as $\sqrt{1-\psi_{q}}$ and given a random sign ensuring that the genetic component of each secondary feature is simulated on the correlation scale. The loadings of the focal trait on the $m_{s}=4$ latent signal factors were obtained as $\sqrt{(1-\psi_{p+1})/m_{s}}$. The focal trait thus loads equally on all four latent signal factors.
The genetic components and corresponding correlation matrix can now be obtained as $\mathbf{G}=\bm{\Xi}\mathbf{\Lambda}^{\top}+\bm{\Gamma}$ and $\mathbf{R}^{g}=\mathbf{\Lambda}\mathbf{\Phi}\mathbf{\Lambda}^{\top}+\mathbf{\Psi}$. The non-genetic or residual components are sampled from a multivariate normal distribution:
(15) $\operatorname{vec}(\mathbf{E})\sim\mathcal{N}_{np}\left(\mathbf{0},\mathbf{R}^{\epsilon}\otimes\mathbf{I}_{n}\right)\text{,}$
where $\mathbf{R}^{\epsilon}$ is a correlation matrix containing random residual correlations (mean of 00 and standard deviation of approximately $0.05$) among the $800$ observed features and between those features and the focal trait. We then finally obtain the phenotypic data:
(16) $\mathbf{Y}=\mathbf{Z}\left(\mathbf{G}\mathbf{S}^{g}\right)+\mathbf{E}\mathbf{S}^{\epsilon}\text{,}$
where the genetic and residual standard deviations in $\mathbf{S}^{g}$ and $\mathbf{S}^{\epsilon}$ are obtained through the heritabilities of the secondary features and focal trait. Using the secondary feature heritabilities $h^{2}(s)\in\{0.5,0.7,0.9\}$ we obtain the genetic and residual standard deviations as $\sigma_{g}(s)=\sqrt{h^{2}(s)}$ and $\sigma_{\epsilon}(s)=\sqrt{1-h^{2}(s)}$, given that we restrict $\sigma^{2}_{g}(s)+\sigma^{2}_{\epsilon}(s)$ to be $1$. We use the focal trait heritabilities $h^{2}(y)\in\{0.1,0.3,0.5,0.7,0.9\}$ to similarly obtain the standard deviations for the focal trait. These heritabilities together with the focal trait unique variance are the three parameters we vary. Figure 1 provides a graphical overview of the simulation setting.
In addition to using random residual correlations as given in (15), we also simulated datasets using a low-rank structure for the residual correlations:
(17) $\operatorname{vec}(\mathbf{E})\sim\mathcal{N}_{np}\left(\mathbf{0},\mathbf{\Lambda}_{\epsilon}\mathbf{\Lambda}_{\epsilon}^{\top}+\mathbf{\Psi}_{\epsilon}\otimes\mathbf{I}_{n}\right)\text{,}$
where $\mathbf{\Lambda}_{\epsilon}\in\mathbb{R}^{(p+1)\times 1}$ contains residual loadings of the secondary features and focal trait on a single factor $\xi_{\epsilon}$ (Figure 7). For each simulated dataset, a random subset of $80$ out of $800$ secondary features load on $\xi_{\epsilon}$. Residual loadings and residual unique variances for these features are simulated in the same way as the loadings and unique variances underlying the genetic correlations. The unique residual variance for the focal trait was equal to the unique genetic variance, i.e., $0.8$, $0.5$, or $0.2$. The $720$ secondary features that have no residual loadings on $\xi_{\epsilon}$ had unique residual variances of 1. A graphical overview of this simulation setting can be found in Figure 7.
@@
Figure 2. Prediction accuracies for several methods in the CV1 and CV2 scenarios as we vary the focal trait communality and heritability at a secondary feature heritability of $0.7$ (random residual correlations). The benchmark predictions use the true simulated factor scores and covariance matrices for multivariate genomic prediction. Further details regarding these simulations can be found in Section 3.1. Results for more secondary feature heritabilities can be found in the Appendix
3.2. Results
The impact of genetic parameters like the focal trait heritability and unique variance is clear (Figure 2). With the exception of the high focal trait unique variance setting, the CV1 accuracy of most multivariate approaches is, at worst, on par with that of the univariate model across the range of focal trait heritabilities. Furthermore, CV1 accuracies of all methods, except the neural networks, converge to more or less the same point ($\rho\approx 0.64$) as $h^{2}(y)$ increases, highlighting that secondary features bring the biggest benefit when the focal trait is hard to measure accurately and thus has a low heritability.
Most methods significantly outperform the univariate model in CV2, provided the quality of the secondary features is high enough. Looking at the setting with the highest quality secondary features, i.e., $\bm{\psi}_{y}=0.2$, improvements with respect to the univariate approach reach approximately $40\%$ to $125\%$ depending on the focal trait heritability and model. Overall, glfBLUP performs best, closely followed by MegaLMM. The difference between MegaLMM and glfBLUP is mainly visible at the lowest focal trait heritability. This might be due to the supervised nature of MegaLMM, where noisy focal trait data is detrimental when included in the modelling. The lsBLUP and siBLUP approaches show similar performance with siBLUP being slightly better. While the CV2 accuracy for the neural networks is higher than its CV1 accuracy, both are significantly lower than any of the other models, including the univariate approach. We provide results for two more secondary feature heritabilities in the Appendix.
We also provide results for varying focal trait heritabilities, focal trait unique variances, and secondary feature heritabilities combined with a low-rank structure for the residual correlations in the Appendix (Figure 6). Simulating data using a low-rank structure for the residual correlations instead of random noise impacts relative model performance in several ways. MegaLMM performance increases slightly, while the impact on glfBLUP’s performance is negligible. This results in an average accuracy for MegaLMM that is now slightly higher than for glfBLUP. Accuracies for lsBLUP decrease significantly while siBLUP performance increases, especially at lower focal trait unique variances. Deep learning performance is unaffected, but remains low.
These results are interesting for several reasons. We have introduced additional complexity by simulating data using a low-rank structure for the residual covariances while keeping the structure underlying the genetics the same. This low rank structure for the residual covariances may appear to be confounded with the low-rank structure for the genetics from a modeling perspective. Intuitively, this additional complexity would result in lower prediction accuracies with models that fail to separate genetic and non-genetic sources of variation seeing the largest reduction in performance. While we observed a drop in accuracy for lsBLUP, both siBLUP and MegaLMM show an increase in prediction accuracy. The siBLUP and MegaLMM models both apply dimension reduction to the phenotypic data, rather than the genetic components. MegaLMM decomposes the phenotypic data on all traits into factors and residuals that are then modeled independently, while siBLUP uses phenotypic covariances between secondary features and the focal trait to construct the selection index. MegaLMM and siBLUP thus seem to use the low-rank residual structure that is present in the phenotypic data as signal, possibly explaining the higher prediction accuracies. The glfBLUP approach is practically unaffected, showing prediction accuracies that are very similar regardless of whether a low-rank or random residual structure is simulated. This indicates that glfBLUP appropriately separates genetic and non-genetic sources of variation before performing the dimensionality reduction as the change in residual structure does not impact the genomic predictions.
4. Analysis of CIMMYT wheat data
In addition to the extensive simulations provided in the previous section we also evaluated the different approaches in terms of accuracy and interpretability using wheat data from Centro Internacional de Mejoramiento de Maíz y Trigo (CIMMYT) (Krause et al., 2019). The following two sections provide details on the dataset, pre-processing, and results.
4.1. Data description and pre-processing
@@
We first resolved the experimental design for each secondary feature and yield using the following SpATS (Spatial Analysis of Field Trials with Splines) model implemented in LMMsolver (Rodríguez-Álvarez et al., 2018; Pérez-Valencia et al., 2022; Boer, 2023):
(18) $y_{(i|j)klmt}=\mu+g_{jt}+t_{kt}+r_{lt}+c_{mt}+f_{t}\left(r,c\right)+\epsilon_{(i|j)klmt}\text{,}$
where $y_{(i|j)klmt}$ is yield or any of the secondary features for replicate $i$ given genotype $j$ in trial $k$ with row number $l$ and column number $m$, measured at timepoint $t$. The overall mean is given by $\mu$ and $g_{jt}$ is the fixed genotypic effect. The random trial effect $t_{kt}$ is assumed to be independently, identically, and normally distributed with mean zero and variance $\sigma_{t}^{2}$: $t_{k}\sim\mathcal{N}(0,\sigma_{t}^{2})$. Similarly, $r_{lt}\sim\mathcal{N}(0,\sigma_{r}^{2})$, $c_{mt}\sim\mathcal{N}(0,\sigma_{c}^{2})$, and $\epsilon_{(i|j)klmt}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$ for the random row and column effects, as well as the residual effect. The two-dimensional smooth surface over row and column positions is defined by $f_{t}\left(r,c\right)$. We then created the pseudo-CRD data for each timepoint by simply adding the estimated genetic effects and residuals from model (18):
$y_{(q|j)t}=\mu+\hat{g}_{jt}+\hat{\epsilon}_{(i|j)klmt}\text{.}$
Note that due to the elimination of the design factors we can replace the indices $i$, $k$, $l$, and $m$ by a single index $q$ that denotes replicate $q$ of genotype $j$. For the secondary features we then fit three-level nested hierarchical growth models through the ten (B5IR) and four (HEAT) timepoints and took fitted values as our final corrected data (Pérez-Valencia et al., 2022). We finally created $250$ datasets for each managed treatment by randomly sampling $12$ trials as the test set and the remaining trials as the training set. We note that by creating training and test sets after pre-processing the data we are violating the independence between training and test set to some extent. Fully separated pre-processing of the training and test data is infeasible, however, as we need the complete data to perform the spatial corrections. Furthermore, full separation would require fitting an infeasibly large number of linear mixed models for pre-processing.
We scaled and centered yield as well as the hyperspectral features separately for the training and test set prior to analyzing the data. We calculated the CV1 prediction accuracy as the correlation between predictions and the BLUEs ($\hat{g}_{j}$) obtained from model (18). For the CV2 prediction accuracies we used the parametric correction described by Runcie and Cheng (2019). In addition to including test set hyperspectral measurements from all timepoints (CV2), we also evaluated the impact of including only measurements from the vegetative phase (CV2VEG), again using the parametric correction. We used a redundancy filtering threshold $\tau=0.95$ for all multivariate approaches as this provides a good trade-off between excluding redundant features, reducing computational complexity, and maintaining enough features to identify latent factors. Details on the methods other than glfBLUP can be found in SM Section 4.